OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 Which of the following systems is not the same as solving the system shown? @johnweldon1993

OpenStudy (anonymous):

can you complete the question

OpenStudy (anonymous):

if you need to solve for x and y, then it shouldnèt be bad

OpenStudy (anonymous):

multiply first equation with -4 and second equation with 3 and add the two to solve for x, once you have that you can find y easily

OpenStudy (johnweldon1993):

yeah...is there more to the question? or are we just solving for x and y again??

OpenStudy (anonymous):

this is the all of question and answer choices 7x - 3y = 4 2x - 4y = 1 Which of the following systems is not the same as solving the system shown? 28x- 12y= 16 and -6x+ 12y= -3 14x- 6y= 4 and -14x+ 28y= 1 -28x+ 12y= -16 and 28x- 56y= 14

OpenStudy (johnweldon1993):

okay there we go so 7x - 3y = 4 2x - 4y = 1 so with answer choice 2... it looks like we multiply the top equation by 2....and the bottom by -7...so lets do that 14x - 6y = 8 and -14 + 28y = -7 hmmm..doesn't look the same as the answer choice right? must not be correct :)

OpenStudy (anonymous):

Solve the following system of equations by the substitution method. 10x + 10y = 1 x = y - 3 What is the value of y?

OpenStudy (johnweldon1993):

substitution....well look you already now what x equals dont you? x equals y - 3 you SUBSTITUTE this into the first equation wherever you see an x so 10(y-3) + 10y = 1 distribute that 10 10y - 30 + 10y = 1 combine those 10y's 20y - 30 = 1 add 30 to both sides 20y = 31 divided both sides by 20 y = 31/20

OpenStudy (anonymous):

Which of the following equations could be the result of using the comparison method to solve the system shown? x + 2y = 6 x - 4y = 8